s=n/2(a+L); for L please show me how to solve this step by step.. thank you

Found 2 solutions by checkley71, edjones:
Answer by checkley71(8403)

S=N/2(A+L)
S=N(A+L)/2 now cross multiply.
N(A+L)=2S
AN+AL=2S
AL=2S-AN
L=(2S-AN)/A ANSWER.

Answer by edjones(7569)

s=n/2(a+L)
2s=n(a+L)
2s=na+nL
2s-na=nL
(2s-na)/n=L
.
Ed